OpenStudy (christinabatitsas):

You know the distance between two cities is 140 miles. What is the distance in kilometers? A. 86.996 km B. 225.298 km C. 225.302 km D. 85.445 km

OpenStudy (drummergirl7):

Can you help me with my question please?

OpenStudy (christinabatitsas):

send me the link

OpenStudy (drummergirl7):

It's in the question thingy I am on moble

OpenStudy (drummergirl7):

Never mind I got it! Bye!!!

OpenStudy (johnweldon1993):

Hint* \(\large 1 \text{ Mile} = 1609.34 \text{ Meters}\) \[\large 1000 \text{ Meters} = 1 \text{ Kilometer}\]

OpenStudy (christinabatitsas):

ok, im so bad at this. i have no idea what to do, do i divide 140miles

OpenStudy (johnweldon1993):

Don't worry about it :) Do it out \[\large (140 \text{ Miles})\times (\frac{1609.34 \text{ Meters}}{1 \text{ Miles}})\times (\frac{1 \text{ Kilometer}}{1000 \text{ Meters}})\]

OpenStudy (christinabatitsas):

is it gonna be C. 225.302 km ?

OpenStudy (johnweldon1993):

Now you can see what cancels \[\large (140 \cancel{\text{ Miles}})\times (\frac{1609.34 \cancel{\text{ Meters}}}{1 \cancel{\text{ Miles}}})\times (\frac{1 \text{ Kilometer}}{1000 \cancel{\text{ Meters}}})\] And ye you got it :)

OpenStudy (christinabatitsas):

i put it in my and got that answer

OpenStudy (christinabatitsas):

thank you so much !

OpenStudy (johnweldon1993):

Not a problem!